### v1

@@ 7 The role of own- and other-firm IVs in the demand estimation
Third, the less concentrated the market, the lower the explanatory power of the two BLP-style instruments. This occurs because, as competition among firms becomes more intense, equilibrium markups converge, which in turn weakens the correlation between markups and instruments. For instance, assuming a simple logit demand, when the market is competitive, the values inside the bracket in (7) are not significantly different across products in the market. In fact, this finding aligns with the observation made by armstrong2016large in the extreme case of $J\rightarrow\infty$. [^24] Specifically, Armstrong (2016) examined cases where markets consist of many single-product firms ($J_{t}=F_{t}$). He demonstrated that as $J_{t}\to\infty$, the explanatory power of BLP-style instruments for equilibrium markups diminishes. Cross-firm or market-level variations in equilibrium markups can help maintain identification power in such cases. See Armstrong (2016) for further details. On the other hand, in a more concentrated market, greater variation in markup sizes reduces the sensitivity of the correlation between markups and instruments to variations from unobserved components.
@@ 8 Conclusion
Correctly assessing industry conduct is essential for establishing antitrust policy and evaluating market efficiency. Existing approaches to testing firm conduct often suffer from reduced testing power due to model misspecification and challenges in estimating demand. In this article, we propose a practical and powerful testing procedure that circumvents these limitations. Our method, built upon the Rivers and Vuong (RV) non-nested model selection framework, simply compares the performance of two BLP-style instrument sets – competition IVs and collusion IVs – in first-stage price regressions and interprets statistically significant results as evidence of either competitive or collusive behavior among firms.
Through extensive Monte Carlo simulations, we evaluate the finite-sample performance of our test statistic across various market conditions, including different levels of market concentration, collusion, and internalization of colluding partners’ profits. The results show that our method is robust to model misspecification, alternative functional forms for instruments, and data limitations, performing comparably to, or better than, existing approaches in detecting collusion across various collusive scenarios. The simplicity of our approach despite its high testing power makes it a practical tool for the preliminary diagnosis of industry conduct. By complementing existing methods, our testing framework provides researchers and regulatory authorities with an efficient and effective way to assess firm behavior and guide antitrust interventions.

### v2

@@ 7 The role of own- and other-firm IVs in demand estimation
Third, the less concentrated the market, the lower the explanatory power of the two BLP-style instruments. This occurs because, as competition among firms intensifies, equilibrium markups converge, which in turn weakens the correlation between markups and instruments. For instance, assuming a simple logit demand, when the market is competitive, the values inside the bracket in equation (8) are not significantly different across products in the market. In fact, this finding aligns with the observation made by armstrong2016large in the extreme case of $J\rightarrow\infty$. [^31] Specifically, Armstrong (2016) examined cases where markets consist of many single-product firms ($J_{t}=F_{t}$). He demonstrated that as $J_{t}\to\infty$, the explanatory power of BLP-style instruments for equilibrium markups diminishes. Cross-firm or market-level variations in equilibrium markups can help maintain identification power in such cases. See Armstrong (2016) for further details. On the other hand, in a more concentrated market, greater variation in markup sizes reduces the sensitivity of the correlation between markups and instruments to fluctuations from unobserved components.
8 Empirical application
In this section, we apply our testing method to study industry conduct in two differentiated product markets in South Korea: the new passenger car market and the instant noodles market. We then examine whether the results align with anecdotal and analytical evidence of industry conduct in each market. Namely, automobile manufacturers owned by the same parent company jointly determine the prices of their car models, while instant noodle prices remain significantly below the collusive level despite suspicions from the Korean Fair Trade Commission (KFTC).
8.1 The new passenger car market
South Korea is one of the largest car markets in the world, with approximately 1.6 million vehicles sold in 2022.323232https://www.statista.com/statistics/265891/vehicles-sales-in-selected-countries/ The two largest firms, Hyundai and Kia, together accounted for nearly 70 percent of total vehicle sales over the past decade. Since both are owned by the Hyundai Motor Group, following Kia’s merger into the group in 1998, pricing is likely coordinated between the two brands. [^33] We use the terms “firm” and “brand” interchangeably. We test this joint profit maximization against own-profit maximization, in which each firm sets prices independently. We also consider three alternative firm conduct scenarios: joint profit maximization among German brands; collusion among domestic automakers and among foreign automakers; and full collusion. In February 2023, the KFTC imposed a total fine of 42.3 billion Won (approximately 33 million US dollars) on the four German automakers – Mercedes-Benz, BMW, Audi, and Volkswagen – for colluding on emission reduction technology applied to their diesel passenger cars. [^34] The KFTC’s ruling is available at https://www.ftc.go.kr/www/selectBbsNttView.do?pageUnit=10&pageIndex=92&searchCnd=all&key=12&bordCd=3&searchCtgry=01,02&nttSn=42758. Here, we investigate whether the German automakers also colluded on prices.
Drawing on province/year/product-level data on prices and attributes for 776 products from 13 brands between 2012 and 2023, [^35] We define a product as a unique combination of nameplate and fuel type. Appendix LABEL:app:_data_appendix provides details on the raw data and sample construction. we conduct pairwise hypothesis tests by running the following first-stage price regressions:
$p_{jt}=\gamma^{h}\cdot x_{jt}+\bm{\theta}^{h}\mathbf{z}_{jt}^{h}+\psi_{j}+\psi_{y}^{\textrm{fuel}}+\psi_{r}+e^{h}_{jt},$ (19)
where $x_{jt}$ represents four exogenous characteristics of product $j$ in market (year-province) $t$: fuel economy (km per 1,000 Won), acceleration (horsepower/weight), size (width×$\times$×length×$\times$×height), and a constant term. The regression model (19) also includes controls for product, year-fuel type (to account for differential trends in government policies regarding fuel types), and province, denoted as $\psi_{j}$, $\psi_{y}^{\textrm{fuel}}$, and $\psi_{r}$.
As for the exogenous instruments under conduct hypothesis $h\in\{H_{0},H_{1}\}$, denoted by $\mathbf{z}_{jt}^{h}$, we use the four exogenous product attributes to construct four summation IVs (sums over own‐firm products based on the effective firm index). In addition, we incorporate a two‐level nesting structure based on the car segment (small/compact, midsize, mid luxury, large/luxury, small SUV, standard SUV, large/luxury SUV) and fuel type (gasoline, diesel, LPG, EV, HEV), [^36] Note that while the nesting order is important for demand estimation, it is irrelevant to our testing procedure. and create 12 additional summation IVs: sums over own‐firm products in the same segment, sums over own‐firm products with the same fuel type, and sums over own‐firm products with the same fuel type in the same segment. [^37] We obtain testing results consistent with those reported in Table 4 when we proceed without any nesting structures and use just the four sums over own‐firm products as IVs.
We report the testing results (clustered by market) in the upper panel of Table 4 where a positive value greater than 1.65 (a negative value less than -1.65) constitutes evidence for the row (column) conduct model. The results indicate that, as expected, joint profit maximization by the parent company is more consistent with the data than own-profit maximization. Moreover, our test rejects all forms of price collusion against joint profit maximization by the parent company. These results, robust to alternative choices of fixed effects and clustering, align with the fact that no antitrust cases have been filed against the car manufacturers in South Korea in recent years, except for the non-price collusion case described above.
Table 4: Testing firm conduct in the two differentiated product markets
[…]
1. Brand ($F=13$) -3.718 1.013 0.335 4.940
2. Parent company ($F_{c}=9$) 3.718 5.404 2.229 7.104
3. German automakers ($F_{c}=7$) -1.013 -5.404 -0.343 4.954
4. Domestic/Foreign ($F_{c}=2$) -0.335 -2.229 0.343 4.168
5. Full collusion ($F_{c}=1$) -4.940 -7.104 -4.954 -4.168
[…]
1. Competition ($F=4$) 18.727 17.546 18.411 2.619
2. Full collusion ($F_{c}=1$) -18.727 -6.324 -5.444 -22.873
3. N-O-P ($F_{c}=2$) -17.546 6.324 3.747 18.316
4. N-O-S ($F_{c}=2$) -18.411 5.444 -3.747 21.808
5. N-P-S ($F_{c}=2$) -2.619 22.873 -18.316 -21.808
Notes: The table reports the results of pairwise hypothesis tests using our testing framework. Results for the new passenger car market are in the upper panel, while those for the instant noodles market are in the bottom panel. The names of the four ramen producers – Nongshim, Ottogi, Paldo, and Samyang – are abbreviated using their initial letters: N, O, P, and S. A positive value greater than 1.65 (a negative value less than -1.65) constitutes evidence for the row (column, respectively) conduct model.
8.2 Instant noodles market
The South Korean instant noodles (ramen) market is highly concentrated, with four firms – Nongshim, Ottogi, Samyang, and Paldo – accounting for over 90 percent of total sales. In March 2012, the KFTC fined them 135 billion won (approximately 120 million US dollars) for price collusion. In particular, Nongshim, the market leader with more than a 50% sales share, was fined 100 billion won for leading the collusion. However, this ruling was overturned by the Supreme Court of Korea in December 2015.383838The rulings of the KFTC and the Supreme Court are available at https://www.ftc.go.kr/www/selectBbsNttView.do?pageUnit=10&pageIndex=507&searchCnd=all&key=12&bordCd=3&searchCtgry=01,02&nttSn=37962 and https://www.scourt.go.kr/supreme/news/NewsViewAction2.work?seqnum=5081&gubun=4&searchOption=&searchWord=, respectively. Given that the KFTC remains suspicious of these firms and has been closely monitoring ramen prices, [^39] For example, the KFTC considered investigating collusion among the ramen manufacturers in 2023, citing that the ramen prices, which increased in the aftermath of the COVID-19 outbreak, did not decrease despite falling flour prices since late 2022: https://www.newsis.com/view/NISX20230623_0002349874. we test various collusion hypotheses, with Nongshim as the collusion leader, against the hypothesis of own-profit maximization. Specifically, we consider full collusion and three cases in which Nongshim coordinates prices with two other firms.
The data used for our test comprise region/year-month/product-level prices and product-level attributes of the 70 best-selling instant noodle products from 2010 to 2019.404040A region is composed of two or more adjacent provinces. See Appendix LABEL:app:_data_appendix for data details. For each of the seven product attributes – cholesterol, calorie, sugar, fat, protein, sodium contents, and a constant term – we calculate the sum over own-firm products (based on the effective firm index) and use them as exogenous instruments. In addition, given that the ramen market is differentiated along two dimensions (hong2023does; kim2024collusion), that is, package (pouch vs. cup) and soup type (red soup, white soup, soupless), [^41] We exclude cold noodle products from the analysis because their demand exhibits seasonality distinct from that of the products included in the analysis. we construct 21 additional IVs: sums over own‐firm products that share the same package, sums over own‐firm products that share the same soup type, and sums over own‐firm products that share both the same package and soup type. Then, we regress the price on product, time (year-month), and region dummy variables, along with the 28 IVs, for each conduct scenario. [^42] Product attributes are absorbed by the product dummies.
According to the testing results (clustered by market) presented in the bottom panel of Table 4, our test rejects all collusion hypotheses considered in favor of the own-profit maximization. [^43] We also consider other collusion scenarios in which Nongshim, along with another firm, coordinate prices, and obtain results in favor of price competition. These results align with the 2015 Supreme Court ruling, as well as the findings of kim2024collusion concluding that the observed markups are too low to support any collusive behavior.
@@ 9 Conclusion
Correctly assessing industry conduct is essential for establishing antitrust policy and evaluating market efficiency. Existing approaches to testing firm conduct often suffer from reduced testing power due to model misspecification and challenges in demand estimation. In this article, we propose a practical and powerful testing procedure that circumvents these limitations. Our method, built upon the Rivers and Vuong (RV) non-nested model selection framework, simply compares the performance of two BLP-style instrument sets – competition IVs and collusion IVs – in first-stage price regressions and interprets statistically significant results as evidence of either competitive or collusive behavior among firms.
Through extensive Monte Carlo simulations, we evaluate the finite-sample performance of our test statistic under various market conditions, characterized by different levels of market concentration, collusion, and internalization of colluding partners’ profits. The results show that our method is robust to model misspecification, alternative functional forms for instruments, and data limitations, performing comparably to, or better than, existing approaches in detecting collusion across various collusive scenarios. The simplicity of our approach despite its high testing power makes it a practical tool for the preliminary diagnosis of industry conduct. By complementing existing methods, our testing framework provides researchers and regulatory authorities with an efficient and effective way to assess firm behavior and guide antitrust interventions.